a beam of particles strikes a wall containing 2x10^29 nuclei per m^3 .each nucleus behaves like a sphere of radius 3x10^-15 m.find the thickness of a wall that exactly half the particles will penetrate without scattering .what thickness would be needed to stop all but one particle in 10^6?

1
Expert's answer
2022-05-11T13:52:29-0400

"d=\frac{\sqrt\frac rn}2=6.1\cdot 10^{-8}~m."